1. Find the radius of curvature (ROC) at (a cos3 θ, a sin3


θ) on the curve x

2/3 +


y

2/3 = a

2/3



2. Find the radius of curvature (ROC) for the curve x = a(cos θ + θ sin θ), y =

a(sin θ − θ cos θ).


1
Expert's answer
2021-12-22T05:27:55-0500

1.

The parametric equations are


"x=a\cos^3 \theta=\dfrac{a}{4}(3\cos \theta+\cos 3\theta),"

"y=a\sin^3 \theta=\dfrac{a}{4}(3\sin \theta-\sin 3\theta),"

"x'=\dfrac{a}{4}(-3\sin \theta-3\sin 3\theta)"

"x''=\dfrac{a}{4}(-3\cos \theta-9\cos 3\theta)"

"y'=\dfrac{a}{4}(3\cos \theta-3\cos 3\theta)"

"y''=\dfrac{a}{4}(-3\sin \theta+9\sin 3\theta)"

"R=\dfrac{((x')^2+(y')^2)^{3/2}}{|x'y''-x''y'|}"

"(x')^2+(y')^2"

"=\dfrac{a^2}{16}(9\sin^2 \theta+9\sin^2 3\theta+18\sin\theta \sin 3\theta)"

"+\dfrac{a^2}{16}(9\cos^2 \theta+9\cos^2 3\theta-18\cos\theta \cos 3\theta)"

"=\dfrac{9a^2}{8}(1-\cos4\theta)=\dfrac{9a^2}{4}\sin^22\theta"


"((x')^2+(y')^2)^{3/2}=\dfrac{27a^3}{8}\sin^32\theta"

"x'y''-x''y'"

"=\dfrac{a^2}{16}(9\sin^2 \theta-27\sin^2 3\theta-18\sin\theta \sin 3\theta)"

"-\dfrac{a^2}{16}(-9\cos^2 \theta+27\cos^2 3\theta-18\cos\theta \cos 3\theta)"

"=-\dfrac{9a^2}{8}(1-\cos4\theta)=-\dfrac{9a^2}{4}\sin^22\theta"

"R=\dfrac{\dfrac{27a^3}{8}\sin^32\theta}{|-\dfrac{9a^2}{4}\sin^22\theta|}"

"=\dfrac{3a}{2}\sin2\theta=3a\sin\theta \cos\theta"

"R=3a\sin\theta \cos\theta"

2.


"x=a(\cos \theta+\theta \sin \theta)"

"y=a(\sin \theta-\theta \cos \theta)"

"x'=a(-\sin \theta+\sin \theta+\theta\cos \theta)=a(\theta \cos \theta)"

"x''=a(\cos \theta-\theta \sin \theta )"

"y'=a(\cos \theta-\cos \theta+\theta \sin \theta)=a(\theta \sin \theta)"

"y''=a(\sin \theta+\theta \cos \theta )"

"R=\dfrac{((x')^2+(y')^2)^{3/2}}{|x'y''-x''y'|}"

"(x')^2+(y')^2=a^2\theta^2(\cos62\theta+\sin^2\theta)=a^2\theta^2"

"((x')^2+(y')^2)^{3/2}=a^3\theta^3"

"x'y''-x''y'=a^2(\theta\sin\theta \cos\theta+\theta^2\cos^2\theta)"

"-a^2(\theta\sin\theta \cos\theta-\theta^2\sin^2\theta)=a^2\theta^2"

"R=\dfrac{a^3\theta^3}{|a^2\theta^2|}=a\theta"